OpenStudy (anonymous):

Please help!!! how do you multiphy x^2/5*^2/9?

OpenStudy (anonymous):

http://www.algebrahelp.com/calculators/equation/

OpenStudy (anonymous):

plug it in

OpenStudy (anonymous):

itll give you your variable

OpenStudy (mathmate):

Do you mean \((\frac{x^2}{5})^2\div9\) ?

OpenStudy (mathmate):

or \(\large (\frac{x^2}{5})\times \frac{2}{9}\)?

OpenStudy (anonymous):

No sorry Mathmate, it's x ^3/5 divided by x^1/4

OpenStudy (anonymous):

Did that make more sense?

OpenStudy (anonymous):

It's the same problem just different numbers